Pedregon Corporation has provided the following information:

Cost per Unit Cost per Period
Direct materials $ 7.00
Direct labor $ 3.60
Variable manufacturing overhead $ 1.25
Fixed manufacturing overhead $ 10,800
Sales commissions $ 0.45
Variable administrative expense $ 0.55
Fixed selling and administrative expense $ 3,200

If 3,000 units are sold, the total variable cost is closest to:
Answer #1

Total Variable cost per unit=Direct materials+Direct labor+Variable manufacturing overhead+Sales commissions+Variable administrative expense

=(7+3.6+1.25+0.45+0.55)=$12.85 per unit

Hence total variable costs would be

=12.85*3000 units

=$38550